OpenStudy (anonymous):

Determine the representation of each of the numbers into the indicated base or notation: 346 to base five

jimthompson5910 (jim_thompson5910):

Similar Example 380 to base five You divide 380 by 5 to get 380/5 = 76 remainder 0 then you take the quotient above (76) and divide it by 5 to get 76/5 = 15 remainder 1 Repeat the last step and you do this until you get a quotient of 0 15/5 = 3 remainder 0 3/5 = 0 remainder 3 Read the remainders in reverse order (from bottom to top) to get 3010 So 380 base 10 = 3010 base 5 Note: this is NOT your answer since it's only an example